determine the values of r for which the given differential equation has solutions of the form y = e^rt

11.y′′′\:−\:3y′′\:+\:2y′\:=\:0


1
Expert's answer
2021-09-06T13:11:30-0400

"\text{The auxilliary equation for the given differential equation is}\\
y^3-3y^2+2y=0\\
\text{Next, we factorise the auxillary equation, since $y-1$ is a factor, we have that}\\
(y^2 + ay + b)(y-1) = y^3 -3y^2 +2y\\
=y^3 +y^2(a-1) +y(b-a) = y^3 -3y^2 +2y\\
\text{Comparing co-efficients, we have that}\\
a-1 = -3 \therefore a=-2\\
b+2 = 2 \therefore b = 0\\
\implies (y^2-2y)(y-1)=0\\
=y(y-1)(y-2)=0\\
\text{Hence $y =0$, $y = 1$ and $y=2$, we know that the solution of the differential equation is}\\
x = ce^{yt}\\
\therefore \text{the values of r are 0, 1 and 2}"